Iodination of Acetone

1.What is reaction kinetics?
2. What variables can affect the rate of a reaction?
3. Given the rate law :    Rate = k [A][B]3      if the concentration of B doubles how will that affect the rate?
4. For a mixture that is : 10mL 4.0M acetone, 10mL 1.0M HCl, 10mL 0.0050M I2, and 20mL water, what is the molarity of each substance IN the mixture?
Answer #1

1) Reaction Kinetics is the thing that provide the information of rate of reaction i.e. how slowly or fast a reaction is occurring.

2) Factors affecting rate of reactions are below:

A) Catalyst

B) Temperature

C) Concentration

D) Number of effective collisions.

3) r = k[A][B]^3

By doubling [B], rate will become

r = k[A][2B]^3 =8* k[A][B]^3

Thus rate is increased by 8 times ..... Answer

4) Moles of Acetone = Molarity * volume in liters = 4*10/1000 = 0.04

Moles of HCl = 1*10/1000 = 0.01

Moles of I2 = 0.0050*10/1000 = 0.00005

Total volume = V = 10+10+10+20 = 50 ml = 0.05 liters

Now,

Molarity of Acetone = 0.04 (moles)/0.05 (total volume) = 0.8 M

Molarity of HCl = 0.01/0.05 = 0.2 M

Molarity of I2 = 0.00005/0.05 = 0.001 M